# Differentially Private Data Generation with Missing Data

### v2

@@ 4.2. Adaptive Recourse Approach
GAN-based adaptive recourse. In non-private literature, several approaches use the GAN framework to deal with missing data (Yoon et al., 2018; Li et al., 2019; Luo et al., 2019; Xu et al., 2020). We choose to privatize an approach called MisGAN (Li et al., 2019), which allows us to simultaneously demonstrate the power of learning the data distribution and the missingness pattern for GAN-based algorithms. We call its DP version, DP-MisGAN, as shown in Algorithm 1. We first describe the high-level architecture of MisGAN and DP-MisGAN, and then we highlight the enhancement in this approach.
Algorithm Overview: MisGAN/DP-MisGAN trains two generator-discriminator pairs — one for learning the data distribution and the other for learning the missingness pattern. The training spans $E$ epochs, with each epoch sampling $|D|/B$ sized subsets from the dataset $D$ without replacement (Line 4). Each subset $S_{t}$ is processed with real data $x_{data}$ (Line 6) and its corresponding missing mask $x_{mask}$(Line 7). The missing mask $x_{mask}$, computed from the missing indicator matrix $M$, marks where data is missing as $1$ and 00 otherwise. Missing values in real data are replaced with 00s (Line 8). Two fake examples $y_{data}$ and $y_{mask}$ are generated by passing random Gaussian noise through data and mask generators (Line 9). These generators are updated using gradient descent, with discriminators learning true distributions in one phase (Line 10) and generators updating within specified $T_{G}$ intervals in the second phase (Line 11). In each generator interval, two fake samples $y_{data}$ and $y_{mask}$ are again generated similarly from the two generators (Line 12) and gradients from the discriminator is computed (Line 13). Note that the non-private MisGAN uses these gradients directly to update the parameters of the two generators (Line 16) and releases both generators in the end. Finally, after the training is completed for $E$ epochs, the discriminators are thrown away, and the privately learned data generator is used to sample synthetic data (Line 20) .
@@
8:        Fill missing values in $x_{data}$ with 00
@@
Highlights: In our DP-MisGAN, we clip and then noise the gradients learned for the generators using the sampled Gaussian mechanism (SGM) (Mironov et al., 2019) to ensure privacy (Lines 14-15). This step is different from prior efforts for DPCTGAN (Fang et al., 2022) and DPautoGAN (Tantipongpipat et al., 2021), which privatize non-private optimizer of the discriminators using DPSGD (Williams and McSherry, 2010; Bassily et al., 2014; Song et al., 2013; Abadi et al., 2016). This divergence in technique leads to notable difference in the algorithm’s utility: while the discriminators become noisy due to the introduction of noise, the generator’s gradient calculation, which relies on discriminator weights (Line 13), is also affected. We observe that we only need to publish the generators, not the discriminators. Hence, we resort to the gradient sanitization (GS) approach (Chen et al., 2020) to perturb only the gradients of the generators (Lines 14-15), without affecting the utility of the discriminator. In addition, the GS approach also allows us to skip the hyperparameter tuning for the gradient clip $C$, which can be vastly detrimental if set wrong (Mohapatra et al., 2021). Unlike the standard gradient clipping to bound the sensitivity of the gradient norm by $C$, i.e., $g/max(1,\|g\|_{2}/C)$, we clip the gradient by $g/max(1,\|g\|_{2})$ (Line 15) by considering a Wasserstein-Gan (WGAN) (Arjovsky et al., 2017) framework with an additional gradient penalty term in the loss function of the discriminator that enforces the $\ell_{2}$-norm of the discriminator gradients to be naturally close to $1$. A tight bound for the privacy loss of the training procedure can be calculated using Rényi-DP (RDP).
Definition 4.5 (Rényi-DP (Mironov, 2017)).
@@
The post-processing and composability properties of DP also apply to RDP. Specifically, if a sequence of adaptive mechanisms $\mathcal{M}_{1}$, $\mathcal{M}_{2}$, $\cdots$, $\mathcal{M}_{k}$ satisfy ($\alpha,\epsilon_{1}$)-, ($\alpha,\epsilon_{2}$)-, $\cdots$, ($\alpha,\epsilon_{k}$)-RDP, then the composite privacy loss is ($\alpha,\sum_{i=1}^{k}\epsilon_{i}$)-RDP. The RDP privacy loss of the sampled Gaussian mechanism (SGM) (Mironov et al., 2019) is given using Lemma 4.6.
@@
By the tail bound property of RDP (Mironov, 2017), we can convert the RDP cost of DP-MisGAN to $(\epsilon,\delta)$-DP, where $\epsilon$ is computed by
@@
for a given $\delta$. In practice, given the parameters $\epsilon,\delta,B,T,T_{G}$, the $\sigma$ is calculated using Theorem 4.7 and the order $\alpha$ is usually searched within a range of user input values (Waites, 2019).
@@
13:     Set negative values to 00 and normalize
@@
Partial marginal observation-based adaptive recourse. This approach can be applied to algorithms that use low dimensional marginal queries (Zhang et al., 2014; McKenna et al., 2019; Qardaji et al., 2014; Xiao et al., 2011; McKenna et al., 2022; Koller and Friedman, 2009; Ping et al., 2017; Li et al., 2014). Instead of discarding all the partially missing rows, only the rows with missing cells in the queried attributes are removed. Such a strategy is most helpful when only a subset of attributes have missing data. For example, with MAR missing mechanism, partial marginal observation can be learned from all the non-missing columns.
Algorithm overview: We extend PrivBayes (Zhang et al., 2014) using this strategy and call it PrivBayes enhanced or PrivBayesE in short (Algorithm 2). Both PrivBayes and PrivBayesE learn Bayesian networks of degree $k$ to know the correlated columns. The network $B$ is initialized by adding the first attribute in the attribute list $\mathcal{A}$. The vertices that been discovered so far are stored in the list $V$ (Lines 1-2). Next the algorithm loops over each attribute in $\mathcal{A}$ (Line 3) and generates $|V|$ choose $k$ sets appended with every attribute seen so far $\mathcal{A}\setminus V$ and stores in a list $\Omega$ (Lines 4-5). The mutual information values for each pair in $\Omega$ is computed and the best one chosen using exponential mechanism using privacy budget $\epsilon_{1}/\mathcal{A}$ is added to the network $B$ (Lines 6-7). PrivBayes/PrivBayesE then generate a synthetic dataset $D^{*}$ using this network $B$ in sequence (Line 10). Marginals are computed for each attribute $X_{i}$ with its most correlated attributes $\Psi_{i}$ using privacy budget $\epsilon/2$ and added to $D^{*}$ (Line 12-14).
@@
Algorithm overview: We extend Kamino (Ge et al., 2021) using this strategy and call in Kamino impute or KaminoI in short (Algorithm  3). Kamino/KaminoI starts with deciding a sequence of attributes based on a given denial constraints $\Psi$ (Line 1). The distribution of the first attribute in the computed sequence is learnt using all the non-missing cells (Line 3). This computed distribution is noised (Line 4) and values are sampled to populate the synthetic dataset (Line 5). For each new attribute $Y$ in the sequence, Kamino/KaminoI learns a private intermediate model which uses all previously visited attributes $X$ to predict the new attribute $Y$ (Lines 7-8). This intermediate model is used to generate the values for the attribute $Y$ in the synthetic data given the sampled values for $X$ (Line 10).
@@
Flexibility: This enhancement is applicable to any method that iterates over the columns of the dataset. Such an algorithmic architecture allows for value imputation as learning progresses. Each time a model predicts the next attribute, it can also impute missing values using the same model, incurring no additional privacy costs. This strategy is particularly effective when missing data correlates with other attributes in the dataset, such as in the missing at random (MAR) scenario. While there’s only one known private approach employing this strategy (Ge et al., 2021), several non-private approaches exist (Xu and Veeramachaneni, 2018; Rivers et al., [n.d.]). Although PrivBayesE can use learned distributions to impute missing values for all visited attributes (Line 12 in Algoirthm 2), it only utilizes low-way marginals compared to KaminoI’s larger models, which benefit more from imputation. Hence, we did not include this imputation step in PrivBayesE.
@@ 5.2. Privacy Amplification Due To MCAR
Missing completely at random (MCAR) enforces an independent probability of having missing rows for each attribute in the dataset. We use these probabilities to tighten the privacy bounds for ground truth data when the missing mechanism is MCAR. The technique we developed is inspired by the seminal work of privacy amplification due to sampling (Balle et al., 2018). The premise of privacy amplification by subsampling is that we run a DP algorithm on some random subset of the data (e.g., sampled Gaussian mechanism, DP-SGD). The subset introduces additional uncertainty, which benefits privacy. Privacy amplification due to subsampling has been shown to work for many sampling methods (e.g., Poisson sampling, sampling with/without replacement) and for neighboring datasets which may differ with replacement or substitution. Privacy amplification by subsampling theorem 5.3 makes this intuition precise.
Theorem 5.3 (Sampling Amplification Theorem (Balle et al., 2018; Steinke, 2022)).
@@
Use case 1: Privacy amplification for complete row only approach. Here we show how to apply Proposition 5.4 to all complete row only approaches (PrivBayes, Kamino, and GAN-based approaches). As these approaches take as input all attributes, the probability of seeing a row without missing values is $\prod_{i=1}^{i=k}(1-\phi_{i})$. The following example illustrates how this probability can be used to obtain a tighter privacy bound for the ground truth data.
@@
Consider the incomplete dataset from Figure 3. Let’s assume that the missing data comes from an MCAR mechanism where the missing probabilities are $\phi_{State}=\frac{1}{4},\phi_{Occupation}=0,\phi_{Gender}=\frac{1}{4},\phi_{Income}=\frac{1}{4}$ Given $4$ DP sub-algorithms $M_{1},M_{2},M_{3},M_{4}$ that each offer DP guarantee to the incomplete dataset $D$ at budget $\frac{\epsilon}{4}$. $M_{1}$ computes the marginals of the complete rows over attribute ¡State¿, $M_{2}$ over ¡Occupation¿, $M_{3}$ over ¡Gender¿ and $M_{4}$ over ¡Gender, Income¿. As all sub-algorithms take as input only the complete rows, using Proposition 5.4, the amplification is $\prod_{i}(1-\phi_{i})=0.421$ and using Theorem 5.6 the final privacy is $\bar{\epsilon}=4*0.421\frac{\epsilon}{4}=0.421\epsilon$.
@@
Given DP mechanisms $M_{1},\ldots,M_{j}$ for computing marginals over attribute sets $\mathcal{A}_{1},\ldots,\mathcal{A}_{j}$, a partition of these mechanisms $P=\{G_{1},\ldots G_{i}\}$ is considered valid if it satisfies these conditions: (1) All mechanisms in the same partition are amplified with the same set of amplification attribute set and with the same amplification factors; and (2) The amplification attribute sets of all partitions are disjoint.
The privacy cost for a valid partition is $\bar{\epsilon}=\sum_{G_{l}\in P}p_{\mathcal{A}_{G_{l}}}\sum_{M_{j}\in G_{l}}\epsilon_{j}$, where $\mathcal{A}_{G_{l}}$ is the amplification attribute set for the mechanisms grouped into $G_{l}$ and $p_{\mathcal{A}_{G_{l}}}$ is the corresponding amplification factor.
@@
0:  Marginals $\mathcal{S}=\{S_{1},\ldots,S_{j}\}$ over attributes $\mathcal{A}$, DP mechanisms =$\mathcal{M}=\{(M_{1},\epsilon_{1}),\ldots,(M_{j},\epsilon_{j})\}$ for $\mathcal{S}$, Missing probabilities $\Phi=\{\Phi_{A}|A\in\mathcal{A}\}$ for MCAR
@@
5:     for each $S_{i}\in\mathcal{S}$ with its corresponding ($(M_{i},\epsilon_{i})$do
@@
7:        Find the best amplification attribute set for $M_{i}$:$\mathcal{A}_{l*}\leftarrow argmin_{\mathcal{A}_{l}\in P_{\mathcal{A}}\wedge\mathcal{A}_{l}\subseteq S_{i}}c_{\mathcal{A}_{l}}$
@@
Discussion on MAR and MNAR. MAR and MNAR mechanisms model probabilities of missing data conditioned upon non-missing values from other columns (MAR) or missing values from the same column in the dataset (MNAR). Unfortunately, verifying the presence of MAR and MNAR missing types in the dataset lacks a foolproof method, as it may rely on observed and unobserved variables and their interactions (as discussed in Section 2). Due to the conditional nature of these missing types, achieving privacy amplification with MAR or MNAR mechanisms is challenging. Each row’s probability of having missing values may vary due to conditional dependence on dataset values. While maximum probabilities or noisy upper bounds can be used for privacy amplification if these probabilities are known or privately learned, they are often associated with multiple patterns and complex relationships. For instance, in the simple example depicted in Figure 3, rows with the highest income value (80k) are missing. Here, the probability of a row having a missing value depends directly on the probability of a row having an income of 80k. To calculate missing probabilities associated with MNAR, not only the highest value in the column but also its probability of occurrence needs to be determined. Without distribution assumptions, such calculations cannot be achieved with DP (Bun et al., 2015). Amplification extensions for MAR and MNAR remain a topic for future research.

### v3

@@ 4.2. Adaptive Recourse Approach
GAN-based adaptive recourse. In non-private literature, several approaches use the GAN framework to deal with missing data (yoon2018gain; li2019misgan; luo2019e2gan; xu2020scigans). We choose to privatize an approach called MisGAN (li2019misgan), which allows us to simultaneously demonstrate the power of learning the data distribution and the missingness pattern for GAN-based algorithms. We call its DP version, DP-MisGAN, as shown in Algorithm 1. We first describe the high-level architecture of MisGAN and DP-MisGAN, and then we highlight the enhancement in this approach.
Algorithm Overview: MisGAN/DP-MisGAN trains two generator-discriminator pairs — one for learning the data distribution and the other for learning the missingness pattern. The training spans $E$ epochs, with each epoch sampling $|D|/B$ sized subsets from the dataset $D$ without replacement (Line 4). Each subset $S_{t}$ is processed with real data $x_{data}$ (Line 6) and its corresponding missing mask $x_{mask}$(Line 7). The missing mask $x_{mask}$, computed from the missing indicator matrix $M$, marks where data is missing as $1$ and 0 otherwise. Missing values in real data are replaced with 0s (Line 8). Two fake examples $y_{data}$ and $y_{mask}$ are generated by passing random Gaussian noise through data and mask generators (Line 9). These generators are updated using gradient descent, with discriminators learning true distributions in one phase (Line 10) and generators updating within specified $T_{G}$ intervals in the second phase (Line 11). In each generator interval, two fake samples $y_{data}$ and $y_{mask}$ are again generated similarly from the two generators (Line 12) and gradients from the discriminator is computed (Line 13). Note that the non-private MisGAN uses these gradients directly to update the parameters of the two generators (Line 16) and releases both generators in the end. Finally, after the training is completed for $E$ epochs, the discriminators are thrown away, and the privately learned data generator is used to sample synthetic data (Line 20) .
@@
8:   Fill missing values in $x_{data}$ with 0
@@
Highlights: In our DP-MisGAN, we clip and then noise the gradients learned for the generators using the sampled Gaussian mechanism (SGM) (DBLP:journals/corr/abs-1908-10530) to ensure privacy (Lines 14-15). This step is different from prior efforts for DPCTGAN (fang2022dp) and DPautoGAN (tantipongpipat2021differentially), which privatize non-private optimizer of the discriminators using DPSGD (WM10; BST14; song2013stochastic; DBLP:conf/ccs/AbadiCGMMT016). This divergence in technique leads to notable difference in the algorithm’s utility: while the discriminators become noisy due to the introduction of noise, the generator’s gradient calculation, which relies on discriminator weights (Line 13), is also affected. We observe that we only need to publish the generators, not the discriminators. Hence, we resort to the gradient sanitization (GS) approach (chen2020gs) to perturb only the gradients of the generators (Lines 14-15), without affecting the utility of the discriminator. In addition, the GS approach also allows us to skip the hyperparameter tuning for the gradient clip $C$, which can be vastly detrimental if set wrong (mohapatra2021role). Unlike the standard gradient clipping to bound the sensitivity of the gradient norm by $C$, i.e., $g/max(1,\|g\|_{2}/C)$, we clip the gradient by $g/max(1,\|g\|_{2})$ (Line 15) by considering a Wasserstein-Gan (WGAN) (arjovsky2017wasserstein) framework with an additional gradient penalty term in the loss function of the discriminator that enforces the $\ell_{2}$-norm of the discriminator gradients to be naturally close to $1$. A tight bound for the privacy loss of the training procedure can be calculated using Rényi-DP (RDP).
Definition 4.5 (Rényi-DP (mironov2017renyi)).
@@
The post-processing and composability properties of DP also apply to RDP. Specifically, if a sequence of adaptive mechanisms $\mathcal{M}_{1}$, $\mathcal{M}_{2}$, $\cdots$, $\mathcal{M}_{k}$ satisfy ($\alpha,\epsilon_{1}$)-, ($\alpha,\epsilon_{2}$)-, $\cdots$, ($\alpha,\epsilon_{k}$)-RDP, then the composite privacy loss is ($\alpha,\sum_{i=1}^{k}\epsilon_{i}$)-RDP. The RDP privacy loss of the sampled Gaussian mechanism (SGM) (DBLP:journals/corr/abs-1908-10530) is given using Lemma 4.6.
@@
By the tail bound property of RDP (mironov2017renyi), we can convert the RDP cost of DP-MisGAN to $(\epsilon,\delta)$-DP, where $\epsilon$ is computed by
@@
for a given $\delta$. In practice, given the parameters $\epsilon,\delta,B,T,T_{G}$, the $\sigma$ is calculated using Theorem 4.7 and the order $\alpha$ is usually searched within a range of user input values (waites2019pyvacy).
@@
13:  Set negative values to 0 and normalize
@@
Partial marginal observation-based adaptive recourse. This approach can be applied to algorithms that use low dimensional marginal queries (DBLP:conf/sigmod/ZhangCPSX14; DBLP:conf/icml/McKennaSM19; DBLP:conf/sigmod/QardajiYL14; DBLP:journals/tkde/XiaoWG11; mckenna2022aim; DBLP:books/daglib/0023091; DBLP:conf/ssdbm/PingSH17; DBLP:journals/pvldb/LiXZJ14). Instead of discarding all the partially missing rows, only the rows with missing cells in the queried attributes are removed. Such a strategy is most helpful when only a subset of attributes have missing data. For example, with MAR missing mechanism, partial marginal observation can be learned from all the non-missing columns.
Algorithm overview: We extend PrivBayes (DBLP:conf/sigmod/ZhangCPSX14) using this strategy and call it PrivBayes enhanced or PrivBayesE in short (Algorithm 2). Both PrivBayes and PrivBayesE learn Bayesian networks of degree $k$ to know the correlated columns. The network $B$ is initialized by adding the first attribute in the attribute list $\mathcal{A}$. The vertices that been discovered so far are stored in the list $V$ (Lines 1-2). Next the algorithm loops over each attribute in $\mathcal{A}$ (Line 3) and generates $|V|$ choose $k$ sets appended with every attribute seen so far $\mathcal{A}\setminus V$ and stores in a list $\Omega$ (Lines 4-5). The mutual information values for each pair in $\Omega$ is computed and the best one chosen using exponential mechanism using privacy budget $\epsilon_{1}/\mathcal{A}$ is added to the network $B$ (Lines 6-7). PrivBayes/PrivBayesE then generate a synthetic dataset $D^{*}$ using this network $B$ in sequence (Line 10). Marginals are computed for each attribute $X_{i}$ with its most correlated attributes $\Psi_{i}$ using privacy budget $\epsilon/2$ and added to $D^{*}$ (Line 12-14).
@@
Algorithm overview: We extend Kamino (ge2021kamino) using this strategy and call in Kamino impute or KaminoI in short (Algorithm  3). Kamino/KaminoI starts with deciding a sequence of attributes based on a given denial constraints $\Psi$ (Line 1). The distribution of the first attribute in the computed sequence is learnt using all the non-missing cells (Line 3). This computed distribution is noised (Line 4) and values are sampled to populate the synthetic dataset (Line 5). For each new attribute $Y$ in the sequence, Kamino/KaminoI learns a private intermediate model which uses all previously visited attributes $X$ to predict the new attribute $Y$ (Lines 7-8). This intermediate model is used to generate the values for the attribute $Y$ in the synthetic data given the sampled values for $X$ (Line 10).
@@
Flexibility: This enhancement is applicable to any method that iterates over the columns of the dataset. Such an algorithmic architecture allows for value imputation as learning progresses. Each time a model predicts the next attribute, it can also impute missing values using the same model, incurring no additional privacy costs. This strategy is particularly effective when missing data correlates with other attributes in the dataset, such as in the missing at random (MAR) scenario. While there’s only one known private approach employing this strategy (ge2021kamino), several non-private approaches exist (xu2018synthesizing; riverssynthetic). Although PrivBayesE can use learned distributions to impute missing values for all visited attributes (Line 12 in Algoirthm 2), it only utilizes low-way marginals compared to KaminoI’s larger models, which benefit more from imputation. Hence, we did not include this imputation step in PrivBayesE.
@@ 5.2. Privacy Amplification Due To MCAR
Missing completely at random (MCAR) enforces an independent probability of having missing rows for each attribute in the dataset. We use these probabilities to tighten the privacy bounds for ground truth data when the missing mechanism is MCAR. The technique we developed is inspired by the seminal work of privacy amplification due to sampling (balle2018privacy). The premise of privacy amplification by subsampling is that we run a DP algorithm on some random subset of the data (e.g., sampled Gaussian mechanism, DP-SGD). The subset introduces additional uncertainty, which benefits privacy. Privacy amplification due to subsampling has been shown to work for many sampling methods (e.g., Poisson sampling, sampling with/without replacement) and for neighboring datasets which may differ with replacement or substitution. Privacy amplification by subsampling theorem 5.3 makes this intuition precise.
Theorem 5.3 (Sampling Amplification Theorem (balle2018privacy; steinke2022composition)).
@@
Use case 1: Privacy amplification for complete row only approach. Here we show how to apply Proposition 5.4 to all complete row only approaches (PrivBayes, Kamino, and GAN-based approaches). As these approaches take as input all attributes, the probability of seeing a row without missing values is $\prod_{i=1}^{i=k}(1-\phi_{i})$. The following example illustrates how this probability can be used
to obtain a tighter privacy bound for the ground truth data.
@@
Consider the incomplete dataset from Figure 3. Let’s assume that the missing data comes from an MCAR mechanism where the missing probabilities are $\phi_{State}=\frac{1}{4},\phi_{Occupation}=0,\phi_{Gender}=\frac{1}{4},\phi_{Income}=\frac{1}{4}$ . Given $4$ DP sub-algorithms $M_{1},M_{2},M_{3},M_{4}$ that each offer DP guarantee to the incomplete dataset $D$ at budget $\frac{\epsilon}{4}$. $M_{1}$ computes the marginals of the complete rows over attribute ¡State¿, $M_{2}$ over ¡Occupation¿, $M_{3}$ over ¡Gender¿ and $M_{4}$ over ¡Gender, Income¿. As all sub-algorithms take as input only the complete rows, using Proposition 5.4, the amplification is $\prod_{i}(1-\phi_{i})=0.421$ and using Theorem 5.6 the final privacy is $\bar{\epsilon}=4*0.421\frac{\epsilon}{4}=0.421\epsilon$.
@@
Given DP mechanisms $M_{1},\ldots,M_{j}$ for computing marginals over attribute sets $\mathcal{A}_{1},\ldots,\mathcal{A}_{j}$, a partition of these mechanisms $P=\{G_{1},\ldots G_{i}\}$ is considered valid if it satisfies these conditions: (1) All mechanisms in the same partition are amplified with the same set of amplification attribute set and with the same amplification factors; and (2) The amplification attribute sets of all partitions are disjoint. The privacy cost for a valid partition is $\bar{\epsilon}=\sum_{G_{l}\in P}p_{\mathcal{A}_{G_{l}}}\sum_{M_{j}\in G_{l}}\epsilon_{j}$, where $\mathcal{A}_{G_{l}}$ is the amplification attribute set for the mechanisms grouped into $G_{l}$ and $p_{\mathcal{A}_{G_{l}}}$ is the corresponding amplification factor.
@@
0: Marginals $\mathcal{S}=\{S_{1},\ldots,S_{j}\}$ over attributes $\mathcal{A}$, DP mechanisms $\mathcal{M}=\{(M_{1},\epsilon_{1}),\ldots,(M_{j},\epsilon_{j})\}$ for $\mathcal{S}$, Missing probabilities $\Phi=\{\Phi_{A}|A\in\mathcal{A}\}$ for MCAR
@@
5:  for each $S_{i}\in\mathcal{S}$ with its corresponding ($(M_{i},\epsilon_{i})$) do
@@
7:   Find the best amplification attribute set for $M_{i}$$\mathcal{A}_{l*}\leftarrow argmin_{\mathcal{A}_{l}\in P_{\mathcal{A}}\wedge\mathcal{A}_{l}\subseteq S_{i}}c_{\mathcal{A}_{l}}$
@@
Discussion on MAR and MNAR. MAR and MNAR mechanisms model probabilities of missing data conditioned upon non-missing values from other columns (MAR) or missing values from the same column in the dataset (MNAR). Unfortunately, verifying the presence of MAR and MNAR missing types in the dataset lacks a foolproof method, as it may rely on observed and unobserved variables and their interactions (as discussed in Section 2). Due to the conditional nature of these missing types, achieving privacy amplification with MAR or MNAR mechanisms is challenging. Each row’s probability of having missing values may vary due to conditional dependence on dataset values. While maximum probabilities or noisy upper bounds can be used for privacy amplification if these probabilities are known or privately learned, they are often associated with multiple patterns and complex relationships. For instance, in the simple example depicted in Figure 3, rows with the highest income value (80k) are missing. Here, the probability of a row having a missing value depends directly on the probability of a row having an income of 80k. To calculate missing probabilities associated with MNAR, not only the highest value in the column but also its probability of occurrence needs to be determined. Without distribution assumptions, such calculations cannot be achieved with DP (bun2015differentially). Amplification extensions for MAR and MNAR remain a topic for future research.